OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Show work please! Thanks! 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (saifoo.khan):

where the heck are medals?

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

super saifoo!

OpenStudy (saifoo.khan):

what do u think can be the answer?

OpenStudy (anonymous):

you need the work?

OpenStudy (anonymous):

yes! i need the work!

OpenStudy (anonymous):

multiply first equation by 3, second one by 4

OpenStudy (anonymous):

\[9x+12y=-6\] \[8x-12y=40\] then add to get \[17x=34\] so \[x=2\]

OpenStudy (anonymous):

then replace x by 2 in either equation to get \[6+4y=-2\] \[4y=-8\] \[y=-2\]